OpenStudy (anonymous):

A 0.1647g of hydrocarbon sample is going through combustion process. It produces 0.5694g of CO2 and 0.0845g of H2O. Determine the percentage of these elements in this hydrocarbon.

OpenStudy (anonymous):

i got the answr

OpenStudy (anonymous):

Can u please explain me? I don't know these concepts but excited to learn.

OpenStudy (anonymous):

aadarsh it is just maths that is there above determining percentage 1 mole of Co2 has how much grams in it and how much grams of carbon?

OpenStudy (anonymous):

1 mol of CO2 = 16 grams 12 grams of carbon, rite?

OpenStudy (anonymous):

CxHy + 3/2 O2------> xCO2 + Y/2H20 take out mole of CO2 as u are provided mass and molecularweight

OpenStudy (anonymous):

mass of co2 in 1 mole? yes 12 g of carbon is there in how many grams of co2? oxygen ATOM has mass no 16 use that

OpenStudy (anonymous):

Sorry, mass of 1 mol CO2 = 12 + 16*2 = 44

OpenStudy (anonymous):

yes ! tahts it u got the answer almost...........so now how much carbon is there in above apecified mass of co2?

OpenStudy (anonymous):

that is in 0.5694 grams? how much carbon?

OpenStudy (anonymous):

12/44 * 0.5694?

OpenStudy (anonymous):

yeah same way find out the mass of oxygen in co2 then mass of hydrogen in h2o then wat u learnt in math to find the percent of an element in a given substance wat do u do?

OpenStudy (anonymous):

Yeah!!!!!!!! Got it!!!!!!!!!

OpenStudy (anonymous):

first wat i learnt was that never to avoid doing something because u think u cant do it..u may never know that u actually know it well....in other words dont underestimate urself (or over estimate) this was a lesson i learnt and wanted to share with u!

OpenStudy (anonymous):

Yeah that's true. But sometimes fear overcomes me. Still learning to cope up wiht that.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

CxHy + 3/2 O2------> xCO2 + Y/2H20 x=0.5694/44 x=0.01(almost) y=0.0845/18 y=0.04 so simplest ratio is CH4 and percentage comp is 12/16 x 100 = 75 for Carbon and 4/16 x 100= 25 for H